OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 45°. If the distance from the base of the ladder to the wall is 12 feet, how long is the ladder?

OpenStudy (anonymous):



OpenStudy (anonymous):

so we want to find x?

OpenStudy (anonymous):

so do we use tanget?

OpenStudy (anonymous):

does that help or should i explain further??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so far, so the equation is : 45 = x/12 ?

OpenStudy (anonymous):

no.....you should use cosine

OpenStudy (anonymous):

whats the correct equation?

OpenStudy (anonymous):

\[\cos 45^{0} = \frac{ 12 }{ x }\]

OpenStudy (anonymous):

can you continue from there??

OpenStudy (anonymous):

so do I do cos 45 times 12?

OpenStudy (anonymous):

I got the first 1?

OpenStudy (anonymous):

And how would you do this one?

OpenStudy (anonymous):

no.....its rather 12/cos 45

OpenStudy (anonymous):

I got the second?

OpenStudy (anonymous):

so if you input that that on your calculator, what do you get??

OpenStudy (anonymous):

16.97

OpenStudy (anonymous):

yes....so that is it

OpenStudy (anonymous):

what about this question Im confused?

OpenStudy (anonymous):

can you please post it in case i can not help, someone else would

OpenStudy (anonymous):

yes